A balloon in its flat and collapsed state is filled with helium gas at constant pressure 101 kPa slowly. If it attains a sphere of radius 0.8 m during the reversible process, estimate the work done by helium.


1
Expert's answer
2021-06-21T05:52:32-0400

Work done by the gas during the filling process   = "\intop v dp"

"= (\frac{4}{3}\pi r^3) \times 101 =21"6.61 kJ